OpenStudy (anonymous):

i cant remember how to solve this problem y^2-8y-4=0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two rots of x are: \(\huge{\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

ok, so like how do i start?

OpenStudy (anonymous):

the equation is 5y^2-8y-4=0

hartnn (hartnn):

so whats a,b,c? can u find those?

OpenStudy (anonymous):

but thats just it, how do i find a,b,c?

hartnn (hartnn):

 now?